Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Fish Tank!

Expert replies
by pkw209 » Fri Apr 16, 2010 12:37 pm
A fish tank contains orange fish and silver fish. After k more orange fish and 2k more silver fish are added the probability of choosing an orange fish at random is 1/3. What was the probability of choosing a silver fish before any more fish were added?

a) 1/4

b) 1/3

c) 1/2

d) 2/3

e) 3/4
Join the discussion
Source: — Problem Solving |

by eaakbari » Fri Apr 16, 2010 1:04 pm
pkw209 wrote:A fish tank contains orange fish and silver fish. After k more orange fish and 2k more silver fish are added the probability of choosing an orange fish at random is 1/3. What was the probability of choosing a silver fish before any more fish were added?

a) 1/4

b) 1/3

c) 1/2

d) 2/3

e) 3/4
IMO D

From the info given, we infer

o+k/(o+s+3k)= 1/3
Cross multiplying and solving
s=2o

Now calculate prob of s/(o+s)

2s/(s/2 + s)

= 2/3

Hence D
Whether you think you can or can't, you're right.
- Henry Ford
Join the discussion

by liferocks » Fri Apr 16, 2010 5:55 pm
Agree with eaakbari ans should be [spoiler]D)2/3[/spoiler]
just want to share my approach

since probability of choosing orange fish is 1/3 let us consider number of orange fish is 100 and total fish is 300

before adding fish ,number of orange fish is 100-k and total number is 300-k

so probability of selecting orange fish is (100-k)/(300-3k) or 1/3
hence probability of selecting silver fish {1-(1/3)} or 2/3
Join the discussion

by hachette » Fri Apr 16, 2010 8:30 pm
The probabilty of choosing orange or silver fish is 1.
If probability of one is given here it is 1/3, the other must be 1 - 1/3 = 2/3.
Join the discussion

by liferocks » Fri Apr 16, 2010 8:50 pm
hachette wrote:The probabilty of choosing orange or silver fish is 1.
If probability of one is given here it is 1/3, the other must be 1 - 1/3 = 2/3.
can you please explain why in the approach you have taken 1/3 as the probability of choosing orange fish BEFORE 1/3 as in question it is mentioned as AFTER . is it something we can infer ?
Join the discussion

by hachette » Fri Apr 16, 2010 9:50 pm
pkw209 wrote:A fish tank contains orange fish and silver fish. After k more orange fish and 2k more silver fish are added the probability of choosing an orange fish at random is 1/3. What was the probability of choosing a silver fish before any more fish were added
After the addition ,
assume no of orange fish = x silver fish is 2x
Prob(O)= x/x+2x = 1/3

Before addition,
Orange fish = x-k Silver fish = 2x-2k = 2(x-k)
Prob (S) = 2(x-k)/{(x-k) + 2 (x-k)} = 2(x-k)/3(x-k)
Join the discussion